y = x2 double integral of the area of the y = 4x - x? region bounded by parabolas Calculate with the help of
Answer #1

y = - (x² - 4x + 4) +7 U y = 4x - x ² = (4-4)=-(x-2) 2 (0-2) = -4-4) T y-9 so Area of the region IR R is SS dA - A = 2. - R H